# DF40: Toward Next-Generation Deepfake Detection

### v1

@@ Abstract
We propose a new comprehensive benchmark to revolutionize the current deepfake detection field to the next generation. Predominantly, existing works identify top-notch detection algorithms and models by adhering to the common practice: training detectors on one specific dataset ($e.g.$, FF++ [67]) and testing them on other prevalent deepfake datasets. This protocol is often regarded as a "golden compass" for navigating SoTA detectors. But can these stand-out "winners" be truly applied to tackle the myriad of realistic and diverse deepfakes lurking in the real world? If not, what underlying factors contribute to this gap? In this work, we found the dataset (both train and test) can be the "primary culprit" due to the following: (1) forgery diversity: Deepfake techniques are commonly referred to as both face forgery (face-swapping and face-reenactment) and entire image synthesis (AIGC, especially face). Most existing datasets only contain partial types of them, with limited forgery methods implemented ($e.g.$, 2 swapping and 2 reenactment methods in FF++); (2) forgery realism: The dominated training dataset, FF++, contains out-of-date forgery techniques from the past five years. "Honing skills" on these forgeries makes it difficult to guarantee effective detection generalization toward nowadays’ SoTA deepfakes; (3) evaluation protocol: Most detection works perform evaluations on one type, $e.g.$, train and test on face-swapping types only, which hinders the development of universal deepfake detectors. To address this dilemma, we construct a highly diverse and large-scale deepfake detection dataset called DF40, which comprises 40 distinct deepfake techniques (10 times larger than FF++). We then conduct comprehensive evaluations using 4 standard evaluation protocols and 7 representative detection methods, resulting in over 2,000 evaluations. Through these evaluations, we provide an extensive analysis from various perspectives, leading to 12 new insightful findings contributing to the field. We also open up 5 valuable yet previously underexplored research questions to inspire future works.
Note: We will release the DF40 dataset with easy-to-use instructions, a unified codebase for train and test, and pre-trained weights of all models on GitHub. For a quick overview of our conclusions, please refer to the takeaways in Sec 9.
@@ 1 Introduction
We are currently in 2024, an explosive era of Artificial Intelligence Generated Content (AIGC), a world where you can effortlessly make anyone say anything at any time, a realm where reality and fiction blend seamlessly, creating a mesmerizing tapestry of digital artistry and potential deception. Amidst this AIGC revolution, the ease of generating deepfakes [^1] Deepfake is a term derived from ”deep” (the use of deep nets) and ”fake” (does not exist in the real world). has become a "double-edged sword." Unfortunately, deepfake is often misused for manipulating one person’s identity (face-swapping) or controlling facial expressions and movements in a portrait (face-reenactment). It can be particularly harmful as it may lead to severe digital crimes and undermine social trust. Therefore, there is an urgent need to develop a reliable system for detecting deepfakes. In a narrow sense, deepfake refers to face forgery, generally including face-swapping and face-reenactment [94]. In a broader sense, it also includes entire image synthesis (not limited to face), widely called AIGCs [81]. Naturally, deepfake detection methods can be broadly classified into two sub-fields: face forgery detection and AIGC detection. Previous works often treat these sub-fields as separate issues, as it is widely believed that they may have distinct fake artifacts: face forgery can be local, with only the facial region manipulated, while AIGC artifacts can be global, involving the entire synthesized image content. For this reason, most detection methods can only be applied to solve one type of these issues.
@@
To illustrate, in face forgery detection, blending-based detectors [11; 46; 70; 45; 95] have achieved SoTA results on existing datasets ($e.g.$, CDF [47]), which involve image blending to create pseudo-fake training fakes. All these methods are based on an assumption that existing face-swapping should share a blending step [^2] Blending the altered face into an existing background image, as initially proposed by [45].. However, actually, recent advancements in face-swapping ($e.g.$, Simswap [13], FaceDancer [66]) directly generate all content (including the background) without blending. Regrettably, this crucial observation may not be well-known in the field, as most existing datasets typically belong to "DF-family" [^3] The term ”DF-family” refers to face forgery techniques that involve creating a mask around the facial region and blending the generated deepfake face back into the background image using that mask. ($e.g.$, CDF [47]). As a result, blending-based detection methods ($e.g.$, [100; 70]) can achieve remarkable generalization results on the existing deepfake datasets ($e.g.$, CDF [47]) due to the presence of blending artifacts but fail to generalize well outside the DF-family datasets (forgeries wo blending). As evidenced in [92], [70] achieves 85% AUC on the CDF dataset (from DF-family), but only 47% in testing SD [64] generated images (outside DF-family). In Sec. 4, we comprehensively re-evaluate it on 31 distinct deepfakes (some are nowadays’ SoTA), showing that the generalization capability of current blending-based methods is still limited. Similar examples can be observed in the AIGC detection field. Recent work [87] proposed detecting diffusion-generated images by measuring the error between an input real image and its reconstruction counterpart using a pre-trained diffusion model. However, this detector is still "diffusion-heavy," meaning it only learns the diffusion-specific feature.
Hence, although many previous efforts have pursued detection generalization, they often remain limited in addressing only a subset of deepfake types ($e.g.$, only face-swapping belongs to the "DF-family") and cannot truly non-trivially be applied to generalize all deepfake types. In a real-world scenario with unpredictability and complexity, creating such a "unified" detector is essential. We discovered previous work [82] attempted to detect all CNN-generated images within one unified model, but its scope is still limited to the entire image synthesis, not including recent popular diffusion models and other deepfake types ($e.g.$, face-swapping). Hence, this issue has persistently been challenging, as most existing datasets either contain limited types (see Tab. 1) or are specific to a single type ($e.g.$, DF-family only).
@@
Therefore, we realize that a diverse and comprehensive dataset is the true key to "unlock" the potential of ideal deepfake detection. To this end, we "jump" from the previous protocol settings ($e.g.$, train and test solely on DF-family) and propose a new comprehensive dataset called DF40. Our main contribution can be generally summarized as three folds: (1) Our DF40 implements 40 different deepfake techniques, including face-swapping, face-reenactment, entire face synthesis, and face editing. We even include the just-released DiT [59], PixArt-$\alpha$ [10], and highly popular software DeepFaceLab [18] and HeyGen [28] to simulate real-world deepfakes; (2) We introduce 4 standard protocols for evaluations and show 8 insightful findings to the field. We also open up new research questions and topics to inspire future research toward next-generation deepfake detection. (3) We release all pre-trained weights from our evaluations, allowing other researchers can directly leverage them to develop their algorithms for improved detection. For example, we release the pre-trained model for CLIP-large (a vanilla binary classifier trained on the proposed DF40 dataset), which achieves an AUC of 98% on DeepForensics [35], 95% on FaceShifter [44], 90% on CDF-v2 [47], and 85% on DFDC [20] at the frame level. This performance outperforms almost all other existing SoTA detectors without relying on previously considered beneficial techniques such as blending methods and other complex constraints.
Furthermore, it is important to note that, without our DF40 dataset, we might only see "the tip of the iceberg" and assume that blending might be "all you need" to solve most deepfake problems. However, our work reveals that the generalization capabilities of blending-based detection methods are limited, and there is still much room for improvement in the deepfake detection domain. It is crucial to develop more robust detection methods that can effectively handle diverse deepfake techniques and data sources, going beyond blending artifacts to address the deepfake problem comprehensively. So in this work, we re-evaluate existing deepfake detectors thoroughly and open up several valuable yet previously underexplored research questions to inspire future works. For now, we believe we have entered the next generation of deepfake detection.
@@ 2 Background
This section briefly introduces the background and its related works, structured around three focal points. First, we briefly introduce deepfake generation techniques, including face-swapping (FS), face-reenactment (FR), entire face synthesis (EFS), and face editing (FE). We consider both academic works and popular software online. Second, we delve into deepfake detection, including face forgery detection and AIGC detection. Last, we review existing deepfake datasets and benchmarks, underscoring the need for diverse and comprehensive datasets for next-generation deepfake detection.
@@
Based on previous survey [54], deepfake techniques can be typically classified into four types: face-swapping (FS), face-reenactment (FR), entire face synthesis (EFS), and face editing (FE). (1) Face-swapping: This paper classifies the face-swapping technique into two domains: DF-family and FS-family. (i) DF-family involves creating a mask around the facial region (some even include the neck [50]) and blending the generated deepfake face back into the background image using that mask. Most existing and famous face forgery datasets, such as FF-DF [67], Celeb-DF [47] and DFDC [20], belong to this line. (ii) FS-family methods represent another significant category in face-swapping deepfake generation. These methods typically involve the use of an identity-background encoder. It disentangles a face image’s identity and background information during encoding. Notably, these methods directly generate all content, even the background. Much recent face-swapping research works [90; 13] and popular software ($e.g.$, Roop [65]) are within this line. (2) Face-reenactment: Generally, this technique can be used to modify source faces, imitating the actions or expressions of another face. Differing from face-swapping, face-reenactment techniques are rarely considered in existing datasets. Two commonly used reenactment-based forgeries are Face2Face [77] and NeuralTextures [76]. These two forgeries are implemented in the FF++ dataset. Face2Face employs pairs of original and target faces, using key facial points to generate varied expressions, while NeuralTexture uses rendered images from a 3D face model to migrate expressions. Although these forgeries achieve more realistic visual synthetic results compared to their face-swapping counterparts in FF++, Due to the amazingly rapid development of the AIGC technologies ($e.g.$, Digital Human), these relatively old-fashioned methods cannot represent the modern’ SoTA reenactment methods. Our DF40 implements 13 face-reenactment methods in total, including the classical animation [72; 73], SoTA’s audio-based driven methods [97; 60], image-based driven methods [86; 31; 7]. We also include the well-known best face generation technique, HeyGen [28], in our dataset for evaluation. (3) Entire Face Synthesis: This technique can be generally treated as "Face AIGC." With the rapid development of AIGCs, this technique has achieved remarkably notable improvement. The two widely used technologies to generate synthesis faces are GAN ($e.g.$, StarGAN [14]) and Diffusion models ($e.g.$, StableDiffusion [64]). (4) Face Editing: This technique aims to modify the facial attributes ($e.g.$, age and gender) of the given face images. Most of these works utilize the latent code of StyleGAN [39] to perform editing during GAN inversion. In our DF40 dataset, we also consider the diffusion-based face editing method [32].
2.2 Deepfake Detection
Since deepfake detection can be broadly classified into face forgery detection and (general) AIGC detection (not limited to face), we will briefly introduce the typical works of each sub-field. (1) Face Forgery Detection. Earlier deepfake detection techniques primarily relied on traditional digital image processing methods. These approaches utilized natural video characteristics such as phoneme-viseme mismatches [4], depth of field [34], color components [43], and head poses [96]. However, with the advent of deep learning, the focus has now shifted towards learning-based detectors due to their superior feature extraction capabilities. Following the recent deepfake benchmark [94], Most existing deepfake detectors can be broadly divided into naive, spatial, and frequency. (i) The naive detector employs CNNs to perform binary classification that distinguishes deepfake content from authentic data. Numerous CNN-based binary classifiers have been proposed, $e.g.$, MesoNet [2] and Xception [67]; (ii) Spatial detector delves deeper into specific representation such as forgery region location [56], disentanglement learning [93], image reconstruction [8], erasing technology [79], image blending [70; 45; 11], $etc$; (iii) The frequency detector addresses this limitation by focusing on the frequency domain for forgery detection. SPSL [48] and SRM [52] utilize phase spectrum analysis and high-frequency noises, respectively. F3Net [61] proposes using learnable filters for adaptive mining of frequency forgery clues using frequency-aware image decomposition. In our work, we further make another new type: blending-based detectors due to the distinct training paradigm. These methods specifically focus on the detection of blending artifacts [46; 45; 11], generating forged images during training in a self-supervised manner to boost detector generalization. These techniques aim to extract more comprehensive and discriminative features by leveraging various aspects of the input data. (2) AIGC Detection. As generative models produce increasingly realistic images, the importance of AIGC detection has grown. In earlier studies, AutoGAN [98] revealed the presence of unique artifacts in GAN-generated images, which can be observed periodically in the images’ frequency spectrum. Wang et al. [82] suggested that a classifier trained on one CNN generator, with suitable training data and data augmentation, can generalize effectively to other unseen generators. However, these approaches have exhibited limited generalization capabilities when applied to images generated by diffusion models. Recently, DIRE [87] employed the error between an input image and its reconstruction using a pre-trained diffusion model to detect fake images. Nonetheless, this method is specifically designed for diffusion models and cannot be directly applied to detecting GAN-based models.
2.3 Existing Deepfake Datasets.
Most earlier public deepfake datasets published before 2021 [41; 96; 21; 47; 67; 17] generate deepfakes using only the face-swapping technique within "DF-family." Also, these datasets contain only single or no more than 4 specific manipulation approaches. Notably, DFDC [20] implements 7 deepfake approaches (most are DF-family face-swapping) and extends the data scale to the million-level. After that, ForgeryNet [27] applies 15 methods to create forgery data with a million-level data scale. However, with the advent of both AIGCs and diffusion-based generators, producing realistic content has become increasingly popular and widely seen on social media. Consequently, it is uncertain whether a detector trained on older forgery methods can generalize to today’s SoTA deepfakes, $e.g.$, diffusion-generated content. In response, many recent works [36; 74] have begun to focus on current generative models. A very recent study, DiffusionDeepfake [6], proposes a diffusion-specific dataset containing two mainstream state-of-the-art diffusion generators: Stable-Diffusion [64] and MidJourney [53]. Furthermore, to detect these AIGC-generated images (not limited to face), GenImage [102] introduces a million-scale dataset created by 8 different generative models (including both GAN and Diffusion). However, it is crucial to note that [6] and [102] are still within only the AIGC scope and may not guarantee successful detection of other deepfakes, $e.g.$, face-swapping. A more comprehensive comparison of existing datasets can be referred to Tab. 1.
@@ 35 WhichisReal [88] GAN based None FFHQ [38] Image Test Only 1,000 Hyper-link
Face Editing (FE) 36 CollabDiff [32] Diffusion based CVPR 2023 FF++ & CDF Image Test Only 48,000+ Hyper-link
37 e4e [78] StyleGAN based SIGGRAPH 2021 FF++ & CDF Image Test Only 48,000+ Hyper-link
@@ 40 StyleCLIP [58] StyleGAN based ICCV 2021 CelebA [51] Image Test Only 2,000 Hyper-link
3 DF40 Benchmark
3.1 Research Scope and Summary of DF40.
In this work, our scope specifically focuses on face deepfake detection, including face-swapping (FS), face-reenactment (FR), entire face synthesis (EFS), and face editing (FE). Natural image synthesis (such as art) is not within our scope, but we show the potential of using model training on face deepfake to detect these (non-face) AIGC-generated images. DF40 dataset provides $40$ deepfake approaches with $4$ different types: FS, FR, EFS, and FE. For FS and FR, we provide video format data (over 0.1M video clips in total), and image format for EFS and FE (over 1M+ images in total). We also introduce several latest/popular generation techniques ($e.g.$, PixArt-$\alpha$ [10] in 2024) and widely used software ($e.g.$, HeyGen [28], DeepFaceLab [18]). Furthermore, some classical and representative methods are also included, $e.g.$, FOMM [72]. Our dataset shows the advantages in both data diversity and scale (see Tab. 1).
3.2 Original Data Collection and Deepfake Data Generation.
Real Data:
We consider two mainstream and popular face datasets specifically used in the deepfake detection field: FF++ [67] (c23 version) and CDF [47] as our original data. There are two main reasons why we chose these two data domains as the original data for creating fakes: (1) Consistency with previous work: By using these datasets for training and evaluation, we can align with prior studies and assess whether their conclusions and methods are still valid in the context of our new dataset. For instance, many previous works [45; 26] follow the "cross-manipulation" evaluation protocol to test the generalization ability when the forgery method differs while the data domain remains the same (FF++). Therefore, creating fakes of DF40 using FF++ allows us to extend their results from evaluating solely on four limited and outdated forgery methods to more than 30 distinct and latest deepfakes. (2) Impact study of domain and forgery: Prior studies have not considered the scenario where only the data domain changes while the forgery method remains the same. This omission hinders the analysis of the sole impact of the data domain on model generalization. In our DF40 dataset, we use both FF++ and CDF domains to facilitate the research on the influence of both data domain and forgery method, as detailed in our Finding-6. In essence, the choice of these two data domains enables us to build upon previous work and to conduct a more comprehensive analysis of the factors influencing the performance of deepfake detection models. We also provide real data from other existing datasets ($e.g.$, CelebA [51]) to facilitate the unknown domain evaluations (see Tab. 2 for details). We show more details of these original face datasets in the supplementary.
Fake Data:
To guarantee the diversity of deepfake approaches in the proposed DF40, we introduce and implement 40 distinct deepfake techniques, which are listed in Tab. 2. A detailed description of each deepfake method is provided in the supplementary. They are selected according to perspectives of modeling types, conditional sources, forgery effects, and functions. Formally, we denote $x_{t}(i_{t},a_{t},b_{t})$ as the target subject to be manipulated, which possesses attributes ($i$: person identity, $a$: identity-agnostic content, $b$: external attributes) that can uniquely determine itself, while the source $x_{s}(i_{s},a_{s},b_{s})$ is regarded as the conditional media ($c$), driving the target to change either identity or attributes or even both. We use $f_{sw},f_{re},f_{sy},f_{ed}$ to represent four types of deepfakes in our work: FS, FR, EFS, and FE, respectively. Fig. 1 provides some visual examples for each category in DF40, and Fig. 16 generally illustrates them from the method perspective. (i) FS Approach: From Fig. 16(a), FS aims to replace the content of $x_{t}$ with that of $x_{s}$ preserving the identity $i_{s}$. Formally, $\tilde{x}_{t}(\tilde{i}_{s},a_{t},b_{t})$ only swaps identity $i$ from the source $x_{s}$ to the target $x_{t}$, and the identity-agnostic content $a$ are preserved. The swapping process can be expressed as:
$f_{re}(x_{t}(i_{t},a_{t},b_{t})|c_{a})=\tilde{x}_{t}(i_{t},\tilde{a}_{s},b_{t}).$ (2)
$f_{sy}(n)=\tilde{x}_{t}(\tilde{i}_{t},\tilde{a}_{t},\tilde{b}_{t}).$ (3)
$f_{ed}(x_{t}(i_{t},a_{t},b_{t})|c_{b})=\tilde{x}_{t}(i_{t},a_{t},\tilde{b}_{s}),$ (4)
@@ 3 DF40 Benchmark
Table 3: Same data domain, different forgery types (Protocol-1): Cross-manipulation evaluation of different models on FF domain. Within-forgery evaluations are in gray. FS (FF) denotes all FS data within FF domain, with FR (FF), FS (CDF), etc, being similar.
@@
Table 4: Same forgery types, different data domains (Protocol-2): Cross-domain evaluation of different models on CDF domain. Within-forgery evaluations are in gray. FS (FF) denotes all FS data within the FF domain, with FR (FF), FS (CDF), etc, having similar meanings.
@@
Table 5: Different forgery types, different data domains (Protocol-3): toward real-world open-set evaluation of different models on the unknown domain. FS (FF) denotes all FS data within the FF domain, with FR (FF) and EFS (FF) being similar. BI donates using "BlendFake" only to train. Here, we use ❶ to mark the FS method, ❷ for the FR, ❸ for the EFS, and ❹ for the FE.
@@
Figure 3: Heatmaps of One-Verse-All (OvA) evaluation (Protocol-4): Training the baseline ($i.e.$, Xception) on one fake and testing it on other remaining fakes. We show the within-domain evaluation (FF++ domain) and cross-domain evaluation (CDF domain). We also show the performance "drop" from the FF++ to the CDF. Blue donates all FS methods, Green for FR, and Yellow for EFS.

### v2

@@ Abstract
We propose a new comprehensive benchmark to revolutionize the current deepfake detection field to the next generation. Predominantly, existing works identify top-notch detection algorithms and models by adhering to the common practice: training detectors on one specific dataset ($e.g.$, FF++ [62]) and testing them on other prevalent deepfake datasets. This protocol is often regarded as a "golden compass" for navigating SoTA detectors. But can these stand-out "winners" be truly applied to tackle the myriad of realistic and diverse deepfakes lurking in the real world? If not, what underlying factors contribute to this gap? In this work, we found the dataset (both train and test) can be the "primary culprit" due to the following: (1) forgery diversity: Deepfake techniques are commonly referred to as both face forgery (face-swapping and face-reenactment) and entire face synthesis (especially face). Most existing datasets only contain partial types of them, with limited forgery methods implemented ($e.g.$, 2 swapping and 2 reenactment methods in FF++); (2) forgery realism: The dominated training dataset, FF++, contains out-of-date forgery techniques from the past four years. "Honing skills" on these forgeries makes it difficult to guarantee effective detection generalization toward nowadays’ SoTA deepfakes; (3) evaluation protocol: Most detection works perform evaluations on one type, $e.g.$, training and testing on face-swapping types only, which hinders the development of universal deepfake detectors.
To address this dilemma, we construct a highly diverse and large-scale deepfake detection dataset called DF40, which comprises 40 distinct deepfake techniques (10 times larger than FF++). We then conduct comprehensive evaluations using 4 standard evaluation protocols and 8 representative detection methods, resulting in over 2,000 evaluations. Through these evaluations, we provide an extensive analysis from various perspectives, leading to 7 new insightful findings contributing to the field. We also open up 4 valuable yet previously underexplored research questions to inspire future works. We release our dataset, code, and checkpoints at https://github.com/YZY-stack/DF40.
Corresponding Author
@@ 1 Introduction
We are currently in 2024, an explosive era of Artificial Intelligence Generated Content (AIGC), a world where you can effortlessly make anyone say anything at any time, a realm where reality and fiction blend seamlessly, creating a mesmerizing tapestry of digital artistry and potential deception. Amidst this AIGC revolution, the ease of generating deepfakes [^1] Deepfake is a term derived from ”deep” (the use of deep nets) and ”fake” (does not exist). Our scope mainly focuses on face deepfakes but shows the promising in detecting other non-face deepfakes such as arts. has become a "double-edged sword." Unfortunately, deepfake is often misused for manipulating one person’s identity (face-swapping) or controlling facial expressions and movements in a portrait (face-reenactment). It can be particularly harmful as it may lead to severe digital crimes and undermine social trust. Therefore, there is an urgent need to develop a reliable system for detecting deepfakes.
@@
Unfortunately, most existing detection methods can only handle a subset of deepfake types [65; 87]. To illustrate, in face-swapping deepfake detection, blending-based detectors [65; 43] have achieved SoTA results on existing face-swapping datasets ($e.g.$, CDF [42]). But all these methods heavily rely on the assumption that existing face-swapping should share a blending step [^2] Blending the altered (deepfake-generated) face into an existing background image.. However, recent advancements in face-swapping ($e.g.$, Simswap [9], FaceDancer [61]) directly generate all content (including the background) without blending. So "looking for" the specific blending artifacts for detecting the "non-blending" forgeries such as Simswap could be unrealistic. Regrettably, this crucial observation may not be well-known in the field, as most existing face-swapping deepfake datasets [62; 17; 42] involve blending in their fake data. Similarly, many existing face-reenactment generators ($e.g.$, LIA [82]) and all entire-image-synthesis generators ($e.g.$, DDPM [24]) are also without any blending. Hence, developing detectors on these blend-only datasets might limit the model’s potential ability to detect a broad range of deepfake types. In a real-world scenario with unpredictability and complexity, creating such a "unified" detector is essential. However, this issue has persistently been challenging, as most existing datasets either contain limited types or are specific to forgery types (see Tab. 1).
@@
Therefore, we realize that a diverse and comprehensive dataset is the true key to "unlock" the potential of ideal deepfake detection. To this end, we "jump" from the previous protocol settings ($e.g.$, train and test solely on face-swapping fakes) and propose a new comprehensive dataset called DF40. Our main contribution can be generally summarized as two folds: (1) Our DF40 implements 40 different deepfake techniques, including face-swapping, face-reenactment, entire face synthesis, and face editing. We even include the just-released DiT [54], PixArt-$\alpha$ [8], and highly popular software DeepFaceLab [14] and HeyGen [23] to simulate real-world deepfakes; (2) We introduce 4 standard protocols for evaluations and show 8 insightful findings to the field. We also open up new research questions and topics to inspire future research toward next-generation deepfake detection.
@@ 2 Background
Based on previous survey [50], deepfake techniques can be typically classified into four types: face-swapping (FS), face-reenactment (FR), entire face synthesis (EFS), and face editing (FE). (1) Face-swapping: This paper classifies the face-swapping technique into two domains: DF-family and FS-family. (i) DF-family involves creating a mask around the facial region (some even include the neck [46]) and blending the generated deepfake face back into the background image using that mask. Most existing and famous face forgery datasets, such as FF-DF [62], CDF [42] and DFDC [16], belong to this line. (ii) FS-family methods represent another significant category in face-swapping deepfake generation. These methods typically involve the use of an identity-background encoder. It disentangles a face image’s identity and background information during encoding. Notably, these methods directly generate all content, even the background. Many recent face-swapping research works [85; 9] are within this line. (2) Face-reenactment: Generally, this technique can be used to modify source faces, imitating the actions or expressions of another face. Differing from face-swapping, face-reenactment techniques are rarely considered in existing datasets. Two commonly used reenactment-based forgeries are Face2Face [72] and NeuralTextures [71]. These two forgeries are implemented in the FF++ dataset. Due to the amazingly rapid development of the AIGC technologies ($e.g.$, Digital Human), these relatively old-fashioned methods cannot represent the modern’ SoTA reenactment methods. Our DF40 implements 13 face-reenactment methods in total, including the classical animation [67; 68], SoTA’s audio-based driven methods [92; 55], image-based driven methods [82; 26; 5]. We also include the well-known best face generation technique, HeyGen [23], in our dataset for evaluation. (3) Entire Face Synthesis: This technique can be generally treated as "Face AIGC." With the rapid development of AIGCs, this technique has achieved remarkably notable improvement. The two widely used technologies to generate synthesis faces are GAN ($e.g.$, VQGAN [20]) and Diffusion models ($e.g.$, StableDiffusion [59]). (4) Face Editing: This technique aims to modify the facial attributes ($e.g.$, age and gender) of the given face images. Most of these works utilize the latent code of StyleGAN [35] to perform editing during GAN inversion.
@@ 35 WhichisReal [83] GAN based None FFHQ [34] Image Test Only 1,000 Hyper-link
Face Editing (FE) 36 CollabDiff [27] Diffusion based CVPR 2023 CelebA [47] Image Test Only 48,000+ Hyper-link
37 e4e [73] StyleGAN based SIGGRAPH 2021 FF++ & CDF Image Train & Test 48,000+ Hyper-link
@@ 40 StyleCLIP [53] StyleGAN based ICCV 2021 CelebA [47] Image Test Only 2,000 Hyper-link
2.2 Existing Deepfake Datasets.
Most earlier public deepfake datasets published before 2021 [37; 91; 17; 42; 62; 13] generate deepfakes using only the face-swapping technique that involves a blending process. Also, these datasets contain only single or no more than 4 specific manipulation approaches. Notably, DFDC [16] implements 7 deepfake approaches and extends the data scale to the million-level. After that, ForgeryNet [22] applies 15 methods to create forgery data with a million-level data scale. However, with the advent of AIGCs such as diffusion-based generators, producing realistic content has become increasingly popular and widely seen on social media. Consequently, whether a detector trained on older forgery methods can generalize to today’s SoTA deepfakes is uncertain. In response, many recent works [32; 69] have begun to focus on current generative models. A very recent study, DiffusionDeepfake [4], proposes a diffusion-specific dataset containing two mainstream SoTA diffusion generators: Stable-Diffusion [59] and MidJourney [49]. Furthermore, to detect these AIGC-generated images (not limited to face), GenImage [95] introduces a million-scale dataset created by 8 different generative models. However, both [4] and [95] are within only the entire image synthesis scope and may not guarantee successful detection of other deepfakes, $e.g.$, face-swapping. A more comprehensive comparison of existing datasets can be referred to Tab. 1.
3 DF40 Benchmark
Research Scope and Summary of DF40.
In this work, our scope specifically focuses on face deepfake detection, including face-swapping (FS), face-reenactment (FR), entire face synthesis (EFS), and face editing (FE). Natural image synthesis (such as art) is not within our scope, but we show the potential of using model training on face deepfake to detect these AIGC-generated images. DF40 dataset provides $40$ deepfake approaches with $4$ different types: FS, FR, EFS, and FE. For FS and FR, we provide video format data (over 0.1M video clips in total), and image format for EFS and FE (over 1M+ images in total). We also introduce several latest/popular generation techniques ($e.g.$, PixArt-$\alpha$ [8] in 2024) and online software ($e.g.$, HeyGen [23], DeepFaceLab [14; 29]). Furthermore, some classical and representative methods are also included, $e.g.$, FOMM [67]. Our dataset shows the advantages in both data diversity and scale (see Tab. 1).
Original Data Collection and Deepfake Data Generation.
(1) Real Data: We consider two mainstream and popular deepfake datasets: FF++ [62] (c23 version) and CDF [42] as our original data. The rationale behind this selection is that most previous research [40; 65; 90] follows the evaluation protocol of training on FF++ and testing on CDF. By utilizing these datasets for training and evaluation, we can adhere to previous works and verify if their conclusions and methods still hold in the context of our new dataset. We also provide real data from other existing datasets ($e.g.$, CelebA [47]) to facilitate the unknown domain evaluations (see Tab. 2 for details). We show more details of these original face datasets in the Appendix. (2) Fake Data: To guarantee the diversity of deepfake approaches in the proposed DF40, we introduce and implement 40 distinct deepfake techniques, which are listed in Tab. 2. A detailed description of each deepfake method is provided in the Appendix. Formally, we denote $x_{t}(i_{t},a_{t},b_{t})$ as the target subject to be manipulated, which possesses attributes ($i$: person identity, $a$: identity-agnostic content, $b$: external attributes) that can uniquely determine itself, while the source $x_{s}(i_{s},a_{s},b_{s})$ is regarded as the conditional media ($c$), driving the target to change either identity or attributes or even both.
@@ 3 DF40 Benchmark
We use $f_{sw},f_{re},f_{sy},f_{ed}$ to represent FS, FR, EFS, and FE, respectively. Fig. 1 provides some visual examples for each category in DF40, and Fig. 14 generally illustrates them from the method perspective. (i) Face-Swapping (FS): From Fig. 14(a), FS aims to replace the content of $x_{t}$ with that of $x_{s}$ preserving the identity $i_{s}$; (ii) Face-Reenactment (FR): From Fig. 14(b), FR on $x_{t}(i_{t},a_{t},b_{t})$ preserves its identity $i_{t}$ but has its intrinsic attributes $a_{t}$, $e.g.$, pose, mouth and expression manipulated by a driven variable $c_{a}$ and forms $\tilde{x}_{t}(i_{t},\tilde{a}_{s},b_{t})$; (iii) Entire Face Synthesis (EFS): EFS generates a entirely synthesis face $\tilde{x}_{t}(\tilde{i}_{t},\tilde{a}_{t},\tilde{b}_{t})$. We fine-tune the generative model $G$ using the real data from FF++ and CDF, and then obtain a face-synthesis model $f_{sy}$. Then, we can synthesize new synthesis faces from noise $n$. The general generation process is in Fig. 14(c); (iv) Face Editing (FE): From Fig. 14(d), FE on $x_{t}(i_{t},a_{t},b_{t})$ has its external attributes $b_{t}$ altered, such as facial hair and age, controlled by conditional source $c_{b}$, to obtain $\tilde{x}_{t}(i_{t},a_{t},\tilde{b}_{s})$.
@@
Table 3: Same data domain, different forgery types (Protocol-1): Cross-forgery evaluation of different models on the FF domain. FS (FF) denotes all FS data within the FF domain, with FR (FF), FS (CDF), etc, being similar. The results of the within-forgery evaluations are in gray (as the validation set only).
@@
Table 4: Same forgery types, different data domains (Protocol-2): Cross-domain evaluation of different models on CDF domain. FS (FF) denotes all FS data within the FF domain, with FR (FF), FS (CDF), etc, having similar meanings. FS (FF) and FS (CDF) represent the same forgery methods used but created on different data domains.
@@
Table 5: Different forgery types, different data domains (Protocol-3): toward real-world open-set evaluation of different models on the unknown domain. FS (FF) denotes all FS data within the FF domain, with FR (FF) and EFS (FF) being similar. Here, we use ❶ to mark the FS method, ❷ for the FR, ❸ for the EFS, and ❹ for the FE.
@@
Figure 3: One-Verse-All (OvA) evaluation (Protocol-4): Training the baseline ($i.e.$, Xception) on one fake and testing it on other remaining fakes. We show the cross-forgery evaluations on both the FF++ domain and CDF domain. We also show the performance "drop" from the FF++ to the CDF. Blue donates all FS methods, Green for FR, and Yellow for EFS. In each heatmap, more "red" indicates higher values; "White" means 0.5 AUC (by chance), and "blue" indicates values below 0.5.